X component of a is twice its y component. If the magnitude of the vector is 5(1.414) and it makes an angle of 135 with z axis then the vector is

Grade:11

1 Answers

Vikas TU
4 years ago
Dear student 
ax = 2ay 
Cosy = az/a = cos135 = -1/sqrt(2)
az = -a/sqrt(2) 
ax^2 + ay^2 = 50 
4ay^2 + ay^2 + 25 = 50 
ay^2 =5 
ay = sqrt(5)
ax = 2sqrt(5)